Suppose that T is an Exponential (λ=1) random variable and f(x)≐⌊x⌋

i.e. is the largest integer not more than x.

Find the cdf and pmf for X≐f(T). What is E[f(T)]?
Answer #1

PMF of X = T) is r+1 e-tdt e-r e-r-1 -e-(1 -e-1): = 0, 1, 2, 3, = x - エ = 0otherwise CDF of X is Fx(z) = P(X-z) = 0 if r < 0 i-0

Take eq, say then 1-qp-1-e-1 then E(X) =〉.rpol! = pol〉 _ = py- 〉 = pq × (1 - )2 p1- e1